At constant temperature, the volume of a sample of gas at 724 mm Hg is 156 ml. What will the volume be if the pressure is dropped to 246 Hg?
1
Expert's answer
2018-05-15T06:14:30-0400
Answer:

At constant T, P1*V1 = P2*V2; this is Boyle's law

Now, V2 = P1*V1 / P2 = 724*156 / 246 = 459 ml